# On relaxing the $N$-Reachability Implicit Requirement in NMPC Design

### v1

@@ 1 Introduction
The closed-loop stability of dynamical systems under Nonlinear Model Predictive Control (NMPC) is by now a mature topics as far as enumerating sufficient conditions is concerned. One of the most systematically invoked sufficient conditions for theoretical stability of NMPC schemes is the so-called $N$-reachability condition (Müller et al., 2014; Köhler et al., 2018; Mayne et al., 2000; Rawlings et al., 2017). This condition stipulates that the targeted/desired state, say $x^{\text{ref}}$, is reachable in less that $N$ steps from any state belonging to some subset $\mathbb{X}$ of initial states of interest. Technically, the $N$-reachability assumption is used in the stability proof to infer the existence, for any initial state $x\in\mathbb{X}$, of a sequence of input actions $\tilde{u}(x):=(u_{0},\dots,u_{N-1})$ that steers the state from $x$ to the desired value $x^{\text{ref}}$. The cost value associated to $\tilde{u}(x)$ hence serves as an upper bound on the optimal cost at the next state and, from this, the stability proof generally follows thanks to the Lassalle’s principle (La Salle, 1966) through some well known sequence of technical arguments.  Obviously, when saturation constraints are present, the satisfaction of this reachability assumption depends on the distance from $x$ to the desired state $x^{\text{ref}}$. This observation is not new and as far as theoretical developments are concerned, it triggered research works on the concept of reference/command-governor (Garone et al., 2017; Klaučo et al., 2017) in the MPC literature. In a nutshell, this amounts at amending the user-defined set-point so that a constraints-compatible modified set-point is computed and sent to the original NMPC controller. Traditionally the latter solution induces an important extra-computational cost due to the new (or extended) optimization problem to solve. This makes the additional scheme eligible to be one of those theoretical components that are generally discarded in real-life implementations leaving to the final user the duty of feeding the NMPC controller with reasonably filtered set-points.  This paper proposes an NMPC formulation involving almost no additional computational cost that smoothly enforces the system to get closer to some reachable steady state that is not explicitly computed and which is smoothly and automatically updated as time goes and as a natural consequence of the formulation so that the system asymptotically reaches a neighborhood of the originally targeted state. This is done mainly by adding a specific terminal cost that incorporates a penalty on the derivative of the terminal state.  The idea of penalizing the final derivative’s amplitude has been recently proposed (Alamir and Pannocchia, 2021) in the context of economic MPC, although in a different framework and using different set of sufficient conditions. In particular, the $N$-reachability condition is required in (Alamir and Pannocchia, 2021) to derive closed-loop stability while in the present contribution, a new set of intuitive sufficient conditions is proposed that enable to replace the $N$-reachability assumption for all pairs $(x_{0},x^{\text{ref}})\in\mathbb{X}^{2}$ by the $N$ reachability, from $x_{0}$ of at least, one steady state with lower stage cost. This suggestion provides a simple scalar weight that can be added to the set of setting parameters of MPC that can be then tuned using the certification scheme such as the one proposed in (Alamir, 2024).  In spite of the fact that the proposed formulation incorporates a contraction condition on the stage cost, it strongly differs from the existing proposed contraction-based NMPC (Alamir, 2017; Polver et al., 2025) in that it does not consider the prediction horizon as a part of the decision variable to be updated at each decision instant. Moreover, it does not require a specific contraction-related storage function to be chosen as the contraction is related to the original stage cost.  Beside the above mentioned contribution, the formulation adopted in the present paper explicitly acknowledges the fact that during the computation time $[k,k+1)$, the control $u_{k}$ cannot be viewed as a decision variable. Instead, it should be considered as a part of an extended state $z_{k}:=(x_{k},u_{k})$ that cannot be modified making the control $u_{k+1}$ to be applied at the next sampling interval $[k+1,k+2]$, the first truly freely available degree of freedom within the future sequence of input actions. This leads to a rather non conventional notation that strongly differs from standard texts on provably stable NMPC formulations.  The paper is organized as follows: In Section 2, some definitions and notation are given and the problem is stated. Section 3 states and discusses the working assumption. Section 4 provides the proof of the main results. Section 5 provides an illustrative use-case assessing the relevance of the proposed ideas. Finally, the paper ends with a conclusion that summarizes the paper’s finding.
@@ 2 Definitions and Notation
where
$\Phi_{\gamma}(\xi):=\gamma^{2}\|f_{c}(\xi)\|^{2}+\gamma\ell(\xi)$ (7)
@@
Given a compact set $\mathbb{Z}$ in the extended state space, the following sequence of compact sets can be defined recursively by $\mathbb{Z}^{(+i)}:=\Bigl{\{}F(z)\ |\ z\in\mathbb{Z}^{(+(i-1))}\Bigr{\}}$ with the initialization $\mathbb{Z}^{(+0)}=\mathbb{Z}$ Notice that $\mathbb{Z}^{(+i)}$ represents the set containing all possible values of the extended state $i$-step ahead starting from $\mathbb{Z}$ under a sequence of $i$ control inputs contained in $\mathbb{U}$. The union of all these sets up to $i=N$ is simply denoted hereafter by $\mathcal{Z}(\mathbb{Z})$, namely: $\mathcal{Z}(\mathbb{Z}):=\bigcup_{i=0}^{N}\mathbb{Z}^{(+i)}$ which is a compact set that contains all possible extended states resulting from an initial state in $\mathbb{Z}$ after a number of steps lower than or equal to $N$.
@@ 3 Working assumptions
The stage cost is positive and is such that $\ell(z)=0$ implies that $z$ is a steady pair, namely $f_{c}(z)=0$.
  It is important to notice that, according to this assumption, the set $\mathbb{Z}_{\eta}$ is supposed to be compact. The next assumption is the major relaxed reachability assumption:
@@
    $\ell(z^{\tilde{\mathbf{u}}(z)}_{N}(z)\leq\alpha\ell(z)$ (13)
when multiple candidates of $\tilde{\mathbf{u}}$ exist that satisfy (12) and (13), the sequence implicitly referred to by $\tilde{\mathbf{u}}(z)$ is the one that minimizes the term $S^{\tilde{\mathbf{u}}}(z)$.
@@
There exists positive $\kappa_{1}$ such that for all $z\in\mathbb{Z}^{\star}_{J_{0}}$, the following inequality holds true:
@@
  The last Assumption we need is rather technical although intuitively sound. It can be informally stated as follows: If a sequence $\tilde{\mathbf{u}}\in\mathbb{U}^{N}$ leads to a sufficiently small neighborhood of a steady state, then there is a small deformation of $\tilde{\mathbf{u}}$ that leads to a rigorously steady state with close stage cost. More formally:
@@
there exists a sequence of control $\check{\mathbf{u}}$ satisfying:
$\|f_{c}(z^{\check{\mathbf{u}}}_{N}(z)\|=0\ \text{and}\ \ell(z^{\check{\mathbf{u}}}_{N})\leq\ell(z^{\tilde{\mathbf{u}}}_{N}(z))+O(\varepsilon)$ (17)
@@
    Assumption 4 is a specific instantiation of the implication (14) that is a consequence of the previous assumptions 1-3. The linear form of the r.h.s does represent an additional assumption that is used for the sake of simplicity of argumentation although it can be made more general at the price of some technicalities in the proof.
@@
    Finally Assumption 5 is a reformulation of the implicit function theorem in a non necessarily differentiable setting. A variant of this assumption has already been discussed in (Alamir and Pannocchia, 2021). For a more detailed argumentation regarding the relevance of this assumption, the reader might refer to Section 3 of (Alamir and Pannocchia, 2021) where his assumption is discussed in details.
@@ 4 Main results
The streamline of the proof is very close to the commonly used ones in similar contributions. Namely, the first result uses the property of the trajectory resulting from the sequence of inputs $\tilde{\mathbf{u}}$ invoked in Assumption 3 in order to characterize some terminal properties of the optimal solution. Then this characterization is used in order to prove that the optimal solution is such that the optimal cost decreases along the closed-loop trajectory with a rate that is proportional to $\ell(z)$.
@@
For all $z\in\mathbb{Z}^{\star}_{J_{0}}$, the optimal solution to $P(z)$ satisfies the following terminal conditions: $\|f_{c}(z^{\star}_{N}(z)\|^{2}$ $\leq\dfrac{1}{\gamma^{2}}\Bigl{[}\gamma\alpha+\kappa_{1}\Bigr{]}\times\ell(z)$ (18a) $\ell(z^{\star}_{N}(z))$ $\leq{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\dfrac{1}{\gamma^{2}}}\Bigl{[}\alpha+\dfrac{\kappa_{1}}{\gamma}\Bigr{]}\times\ell(z)$ (18b)
@@
$=f(z^{\mathbf{u}^{+}}_{N-1}(z^{\star}_{1}(z))-z^{\mathbf{u}^{+}}_{N-1}(z^{\star}_{1}(z))$ (28)
@@
$\|f_{c}\bigl{(}z^{\mathbf{u}^{+}}_{N}(z^{\star}_{1}(z))\bigr{)}\|^{2}$ $\leq\dfrac{1}{\gamma^{2}}\Bigl{[}\gamma\alpha+\kappa_{1}\Bigr{]}\times\max_{z\in\mathbb{Z}^{\star}_{J_{0}}}[\ell(z)]+$
@@
For sufficiently high $\gamma$, for all initial extended state $z_{0}\in\mathbb{Z}^{\star}_{J_{0}}$ the closed-loop trajectory $\{z^{\text{cl}}_{k}\}_{k=0}^{\infty}$ under the MPC feedback defined by (6)-(8) is such that: $J^{\star}(z^{\text{cl}}_{k+1})-J^{\star}(z^{\text{cl}}_{k})\leq-\beta\ell(z^{\text{cl}}_{k})+O(1/\gamma)$ (30) $\lim_{k\rightarrow\infty}\ell(z^{\text{cl}}_{k})=O(1/\gamma)$ (31) where $\beta>0$.
@@
Notice that (30) instantiated at $k=0$ concerns the comparison between the two terms $J^{\star}(z_{0})$ and $J^{\star}(z^{\star}_{1}(z_{0}))$ as $z^{\text{cl}}_{0}=z_{0}$ and $z^{\text{cl}}_{1}=z^{\star}_{1}(z_{0})$.  Let $\gamma_{0}>0$ be a sufficiently high number for the choice $\varepsilon_{0}=1/\gamma_{0}$ to be sufficiently small in the sense of Assumption 5. Then Corollary 1 stipulates that there is a sufficiently higher $\gamma>\gamma_{0}$ making the sequence $\mathbf{u}^{+}$ invoked in Corollary 1 satisfying (16) of Assumption 5 for $\varepsilon_{0}=1/\gamma_{0}$. Therefore, for any such $\gamma$ there exists a sequence of control inputs $\check{\mathbf{u}}$ that, when applied to $z^{\star}_{1}(z_{0})$ leads to:
@@
$\ell(z^{\check{\mathbf{u}}}_{N}(z^{\star}_{1}(z_{0})))$ $\leq\ell(z^{\star}_{N}(z_{0}))+O(1/\gamma)$ (33)
@@
$\leq\Phi_{\gamma}(z^{\star}_{N}(z_{0}))-\gamma^{2}\|f_{c}(z^{\star}_{N}(z_{0}))\|^{2}+O(1/\gamma)$
@@
$\Phi_{\gamma}(z^{\check{\mathbf{u}}}_{N}(z^{\star}_{1}(z_{0})))-\Phi_{\gamma}(z^{\star}_{N}(z_{0}))\leq O(1/\gamma)$ (34)
@@
$S^{\check{\mathbf{u}}}(z^{\star}_{1}(z_{0}))-$ $S^{\mathbf{u}^{\star}(z_{0})}(z_{0})=\ell(z^{\check{\mathbf{u}}}(z^{\star}_{1}(z_{0})))-\ell(z_{0})$
[Because of (33)] $\leq\ell(z^{\star}_{N}(z_{0}))-\ell(z_{0})+O(1/\gamma)$ (35)
@@
$S^{\check{\mathbf{u}}}(z^{\star}_{1}(z_{0}))-S^{\mathbf{u}^{\star}(z_{0})}(z_{0})\leq-(1-\alpha-\frac{\kappa_{1}}{\gamma})\ell(z_{0})+O(1/\gamma)$
Therefore, for sufficiently high $\gamma$, the last inequality can be written in the following form
$S^{\check{\mathbf{u}}}(z^{\star}_{1}(z_{0}))-S^{\mathbf{u}^{\star}(z_{0})}(z_{0})\leq-\beta\ell(z_{0})+O(1/\gamma)$ (37)
@@
$J(\check{\mathbf{u}}\ |\ z^{\star}_{1}(z_{0}))\leq J^{\star}(z_{0})-\beta\ell(z_{0})+O(1/\gamma)$
@@
$J^{\star}(z^{\star}_{1}(z_{0}))\leq J^{\star}(z_{0})-\beta\ell(z_{0})+O(1/\gamma)$
which proves (30) for $k=0$.  Now in order to repeat the arguments for $k>0$, we need to prove that $z^{\star}_{1}(z_{0})$ still belongs to $\mathbb{Z}^{\star}_{J_{0}}$ as long as $\ell(z_{0})$ is not in a $(1/\gamma)$-neighborhood of 00. But this is a direct consequence of the inequality:
$J^{\star}(z^{\star}_{1}(z_{0}))\leq J^{\star}(z_{0})-\beta\ell(z_{0})+O(1/\gamma)$

### v2

@@ 1 Introduction
The closed-loop stability of dynamical systems under Nonlinear Model Predictive Control (NMPC) is by now a mature topics as far as enumerating sufficient conditions is concerned. One of the most systematically invoked sufficient conditions for theoretical stability of NMPC schemes is the so-called $N$-reachability conditionMüller et al. (2014); Köhler et al. (2018); Mayne et al. (2000); Rawlings et al. (2017). This condition stipulates that the targeted/desired state, say $x^{\text{ref}}$, is reachable in less that $N$ steps from any state belonging to some subset $\mathbb{X}$ of initial states of interest. Technically, the $N$-reachability assumption is used in the stability proof to infer the existence, for any initial state $x\in\mathbb{X}$, of a sequence of input actions $\tilde{u}(x):=(u_{0},\dots,u_{N-1})$ that steers the state from $x$ to the desired value $x^{\text{ref}}$. The cost value associated to $\tilde{u}(x)$ hence serves as an upper bound on the optimal cost at the next state and, from this, the stability proof generally follows thanks to the Lassalle’s principle La Salle (1966) through some well known sequence of technical arguments.  Obviously, when saturation constraints are present, the satisfaction of this reachability assumption depends on the distance from $x$ to the desired state $x^{\text{ref}}$. This observation is not new and as far as theoretical developments are concerned, it triggered research works on the concept of reference/command-governor Garone et al. (2017); Klaučo et al. (2017) in the MPC literature. In a nutshell, this amounts at amending the user-defined set-point so that a constraints-compatible modified set-point is computed and sent to the original NMPC controller. Traditionally the latter solution induces an important extra-computational cost due to the new (or extended) optimization problem to solve. This makes the additional scheme eligible to be one of those theoretical components that are generally discarded in real-life implementations leaving to the final user the duty of feeding the NMPC controller with reasonably filtered set-points.  This paper proposes an NMPC formulation involving almost no additional computational cost that smoothly enforces the system to get closer to some reachable steady state that is not explicitly computed and which is smoothly and automatically updated as time goes and as a natural consequence of the formulation so that the system asymptotically reaches a neighbourhood of the originally targeted state. This is done mainly by adding a specific terminal cost that incorporates a penalty on the derivative of the terminal state.  The idea of penalizing the final derivative’s amplitude has been recently proposed Alamir and Pannocchia (2021) in the context of economic MPC, although in a different framework and using different set of sufficient conditions. In particular, the $N$-reachability condition is required in Alamir and Pannocchia (2021) to derive closed-loop stability while in the present contribution, a new set of intuitive sufficient conditions is proposed that enable to replace the $N$-reachability assumption for all pairs $(x_{0},x^{\text{ref}})\in\mathbb{X}^{2}$ by the $N$ reachability, from $x_{0}$ of at least, one steady state with lower stage cost. This suggestion provides a simple scalar weight that can be added to the set of setting parameters of MPC that can be then tuned using the certification scheme such as the one proposed in Alamir (2024). In spite of the fact that the proposed formulation incorporates a contraction condition on the stage cost, it strongly differs from the existing proposed contraction-based NMPC Alamir (2017); Polver et al. (2025) in that it does not consider the prediction horizon as a part of the decision variable to be updated at each decision instant. Moreover, it does not require a specific contraction-related storage function to be chosen as the contraction is related to the original stage cost.   As in almost all provably stable formulations, some of the assumptions are difficult to check theoretically. This makes a last statistical certification step such as the one proposed in Alamir (2024) mandatory to any NMPC formulation. This is the only way to get a solid assessment of the stability, the real-time implementability as well as the constraint satisfaction. The theoretical foundations provided in this paper give rationale to consider this formulation as a candidate for such a certification procedure.  The paper is organized as follows: In Section 2, some definitions and notation are given and the problem is stated. Section 3 states and discusses the working assumption. Section 4 provides the proof of the main results. Section 5 provides an illustrative use-case assessing the relevance of the proposed ideas. Finally, the paper ends with a conclusion that summarizes the paper’s finding.
@@ 2 Definitions and Notation
$\text{where}\quad\Phi_{\gamma}(\xi):=\gamma^{2}\|f_{c}(\xi)\|^{2}+\gamma\ell(\xi)$ (7)
@@
Given a compact set $\mathbb{Z}$ in the extended state space, the following sequence of compact sets can be defined recursively by $\mathbb{Z}^{(+i)}:=\Bigl{\{}F(z)\ |\ z\in\mathbb{Z}^{(+(i-1))}\Bigr{\}}$ with the initialization $\mathbb{Z}^{(+0)}=\mathbb{Z}$. Notice that $\mathbb{Z}^{(+i)}$ represents the set containing all possible values of the extended state $i$-step ahead starting from $\mathbb{Z}$ under a sequence of $i$ control inputs contained in $\mathbb{U}$. The union of all these sets up to $i=N$ is simply denoted hereafter by $\mathcal{Z}(\mathbb{Z})$, namely: $\mathcal{Z}(\mathbb{Z}):=\bigcup_{i=0}^{N}\mathbb{Z}^{(+i)}$ which is a compact set that contains all possible extended states resulting from an initial state in $\mathbb{Z}$ after a number of steps lower than or equal to $N$.
@@ 3 Working assumptions
[…]
The next assumption is the major relaxed reachability assumption:
@@
    $\ell(z^{\tilde{\mathbf{u}}(z)}_{N}(z))\leq\alpha\ell(z)$ (13)
[…]
@@
There exists positive $\kappa_{1}$ such that for all $z\in\mathbb{Z}^{\star}_{J_{0}}$, the following inequality:
@@
[…]
The last Assumption we need is rather technical although intuitively sound. It can be informally stated as follows: If a sequence $\tilde{\mathbf{u}}\in\mathbb{U}^{N}$ leads to a sufficiently small neighbourhood of a steady state, then there is a small deformation of $\tilde{\mathbf{u}}$ that leads to a rigorously steady state with close stage cost. More formally:
@@
there exists a feasible sequence of control $\check{\mathbf{u}}$ satisfying:
$\|f_{c}(z^{\check{\mathbf{u}}}_{N}(z)\|=0\ \text{and}\ \ell(z^{\check{\mathbf{u}}}_{N}(z))\leq\ell(z^{\tilde{\mathbf{u}}}_{N}(z))+O(\varepsilon)$ (17)
@@
    Assumption 4 is a specific instantiation of the implication (14) that is a consequence of the previous assumptions 1-3. The linear form of the r.h.s of (15) encompasses a large set of possibilities for which it is possible to upper bounding the non-linear dependency by the linear form for sufficiently high $\kappa_{1}$.
@@
    Finally Assumption 5 is a reformulation of the implicit function theorem in a non necessarily differentiable setting. A variant of this assumption has already been discussed in Alamir and Pannocchia (2021). This assumption might not be satisfied for a class of systems that are not regular enough.
@@ 4 Main results
The chain of arguments used in the proof is standard. Namely, the first result uses the property of the trajectory resulting from the sequence of inputs $\tilde{\mathbf{u}}$ invoked in Assumption 3 in order to characterize some terminal properties of the optimal solution. Then this characterization is used in order to prove that the optimal solution is such that the optimal cost decreases along the closed-loop trajectory with a rate that is proportional to $\ell(z)$.
@@
For all $z\in\mathbb{Z}^{\star}_{J_{0}}$, the optimal solution to $P(z)$ satisfies the following terminal conditions: $\|f_{c}(z^{\star}_{N}(z)\|^{2}$ $\leq\dfrac{1}{\gamma^{2}}\Bigl{[}\gamma\alpha+\kappa_{1}\Bigr{]}\times\ell(z)$ (18a) $\ell(z^{\star}_{N}(z))$ $\leq\Bigl{[}\alpha+\dfrac{\kappa_{1}}{\gamma}\Bigr{]}\times\ell(z)$ (18b) which also leads to $\|f_{c}(z^{\star}_{N}(z))\|\leq\dfrac{1}{\sqrt{\gamma}}\ell^{1/2}(z^{\star}_{N}(z))$ (18c)
@@
$=f(z^{\mathbf{u}^{+}}_{N-1}(z^{\star}_{1}(z)))-z^{\mathbf{u}^{+}}_{N-1}(z^{\star}_{1}(z))$ (28)
@@
$\|f_{c}\bigl{(}z^{\mathbf{u}^{+}}_{N-1}(z^{\star}_{1}(z))\bigr{)}\|^{2}$ $\leq\dfrac{1}{\gamma^{2}}\Bigl{[}\gamma\alpha+\kappa_{1}\Bigr{]}\times\max_{z\in\mathbb{Z}^{\star}_{J_{0}}}[\ell(z)]+$
@@
For sufficiently high $\gamma$, for all initial extended state $z_{0}\in\mathbb{Z}^{\star}_{J_{0}}$ the closed-loop trajectory $\{z^{\text{cl}}_{k}\}_{k=0}^{\infty}$ under the MPC feedback defined by (6)-(8) is such that: $J^{\star}(z^{\text{cl}}_{k+1})-J^{\star}(z^{\text{cl}}_{k})\lesssim-\beta\ell(z^{\text{cl}}_{k})\ \text{as $\gamma\rightarrow\infty$}$ (30) $\lim_{k\rightarrow\infty}\ell(z^{\text{cl}}_{k})\sim 0\text{as $\gamma\rightarrow\infty$}$ (31) where $\beta>0$.
@@
Notice that (30) instantiated at $k=0$ concerns the comparison between the two terms $J^{\star}(z_{0})$ and $J^{\star}(z^{\star}_{1}(z_{0}))$ as $z^{\text{cl}}_{0}=z_{0}$ and $z^{\text{cl}}_{1}=z^{\star}_{1}(z_{0})$.  Let $\gamma_{0}>0$ be a sufficiently high number for the choice $\varepsilon_{0}=1/\sqrt{\gamma_{0}}\times(\max_{z\in\mathbb{Z}}[\ell(z)])$ to be sufficiently small in the sense of Assumption 5. Then Corollary 1 stipulates that there is a sufficiently higher $\gamma>\gamma_{0}$ making the sequence $\mathbf{u}^{+}$ invoked in Corollary 1 satisfying (16) of Assumption 5 for $\varepsilon_{0}=\dfrac{1}{\sqrt{\gamma}}\ell^{1/2}(z^{\star}_{N}(z))$ by virtue of (18c). This sequence, when applied to $z^{\star}_{1}(z_{0})$ leads to:
@@
$\ell(z^{\check{\mathbf{u}}}_{N}(z^{\star}_{1}(z_{0})))$ $\leq{\underbrace{\ell\Bigl{(}\overbrace{z^{\star}_{N}(z_{0})+O(1/\sqrt{\gamma}}^{z^{\mathbf{u}^{+}}_{N}(z^{\star}_{1}(z_{0})})\Bigr{)}+O(1/\sqrt{\gamma})}_{\sim\ell(z^{\star}_{N}(z_{0}))\ \text{as}\ \gamma\rightarrow\infty}}$ (33)
@@
${\lesssim\Phi_{\gamma}(z^{\star}_{N}(z_{0}))-\gamma^{2}\|f_{c}(z^{\star}_{N}(z_{0}))\|^{2}}$
@@
${\Phi_{\gamma}(z^{\check{\mathbf{u}}}_{N}(z^{\star}_{1}(z_{0})))-\Phi_{\gamma}(z^{\star}_{N}(z_{0}))\lesssim 0\ \text{as $\gamma\rightarrow\infty$}}$ (34)
@@
$S^{\check{\mathbf{u}}}(z^{\star}_{1}(z_{0}))-$ $S^{\mathbf{u}^{\star}(z_{0})}(z_{0})=\ell(z^{\check{\mathbf{u}}}_{N}(z^{\star}_{1}(z_{0})))-\ell(z_{0})$
[Because of (33)] $\lesssim\ell(z^{\star}_{N}(z_{0}))-\ell(z_{0})\ \text{as $\gamma\rightarrow\infty$}$ (35)
@@
$S^{\check{\mathbf{u}}}(z^{\star}_{1}(z_{0}))-S^{\mathbf{u}^{\star}(z_{0})}(z_{0})\lesssim-(1-\alpha)\ell(z_{0})\ \text{as $\gamma\rightarrow\infty$}$
Therefore,
$S^{\check{\mathbf{u}}}(z^{\star}_{1}(z_{0}))-S^{\mathbf{u}^{\star}(z_{0})}(z_{0})\lesssim-\beta\ell(z_{0})\ \text{as $\gamma\rightarrow\infty$}$ (37)
@@
$J(\check{\mathbf{u}}\ |\ z^{\star}_{1}(z_{0}))\lesssim J^{\star}(z_{0})-\beta\ell(z_{0})\text{as $\gamma\rightarrow\infty$}$
@@
$J^{\star}(z^{\star}_{1}(z_{0}))\lesssim J^{\star}(z_{0})-\beta\ell(z_{0})\quad\text{ as $\gamma\rightarrow\infty$}$
which proves (30) for $k=0$.  Now in order to repeat the arguments for $k>0$, we need to prove that $z^{\star}_{1}(z_{0})$ still belongs to $\mathbb{Z}^{\star}_{J_{0}}$ as long as $\ell(z_{0})$ is not in a sufficiently small neighbourhood of 00. But this is a direct consequence of the inequality:
$J^{\star}(z^{\star}_{1}(z_{0}))\lesssim J^{\star}(z_{0})-\beta\ell(z_{0})\quad\text{ as $\gamma\rightarrow\infty$}$